# Armijo Line-search Can Make (Stochastic) Gradient Descent Provably Faster

### v2

@@ C.2 Helper Lemmas
$\mathbb{E}[f(\theta_{T})-f(u)]$ $\leq\frac{\epsilon}{2}+\underbrace{[\nu^{2}+1]}_{:=L}\,\mathbb{E}\frac{\left\|\theta_{T}-u\right\|_{2}^{2}}{2}\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left[\beta^{2}-\frac{C\,T}{4}\right]$
Hence, in order to ensure that $\mathbb{E}[f(\theta_{T})-f(u)]\leq\frac{3\epsilon}{2}$, it is sufficient to set $T$ as follows.
$T$ $\geq\frac{4\,\beta^{2}}{C}=\frac{4}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\,\frac{c\lambda_{1}}{(2c-1)}$ (Substituting the value of $\beta$ and $C$)
$=\frac{4}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\,\frac{3\,c\,L_{1}\,(\nu+1)}{(2c-1)\,(1-c)}$ (Substituting the value of $\lambda_{1}$)
$=\frac{864}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\,\frac{c}{(2c-1)\,(1-c)}$ (Substituting $L_{1}=8$, $\nu=8$)
Hence, with this choice of $T$, we can conclude that,
$\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]|G^{c}\right]$ $\leq\frac{3\epsilon}{2}$
Combining this with Eq. 37, we conclude that
$\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]\right]$ $\leq 3\epsilon\implies\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]\leq 4\,\epsilon$ (Since $f(u)\leq\epsilon$)

### v3

@@ C.2 Helper Lemmas
$\mathbb{E}[f(\theta_{T})-f(u)]$ $\leq\frac{\epsilon}{2}+\underbrace{[\nu^{2}+1]}_{:=L}\,\mathbb{E}\frac{\left\|\theta_{T}-u\right\|_{2}^{2}}{2}\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left[\beta^{2}-C\,T\right]$
$\implies\mathbb{E}[f(\theta_{T})]$ $\leq\frac{3\epsilon}{2}+\frac{L}{2}\,\left[\beta^{2}-C\,T\right]$ (Since $f(u)\leq\epsilon$)
In order to ensure that $\mathbb{E}[f(\theta_{T})]\leq\frac{3\epsilon}{2}$, it is sufficient to set $T$ as:
$T\geq\frac{\beta^{2}}{C}$ $=\frac{4\,n}{\min\{C^{\prime},1\}\,\gamma^{2}}\,\left[\ln\left(\frac{n}{\epsilon^{2}}\right)\right]^{2}$
Since the above calculations were conditioned on $G^{c}$, we can conclude that after $T=\frac{4\,n}{\min\{C^{\prime},1\}\,\gamma^{2}}\,\left[\ln\left(\frac{n}{\epsilon^{2}}\right)\right]^{2}$ iterations, $\mathbb{E}[\inf_{t\in[T]}f(\theta_{T})|G^{c}]\leq\frac{3\epsilon}{2}$. Hence, from Eq. 29
$\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]$ $\leq\frac{5\epsilon}{2}$
Corollary 7.
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon\in\left(0,\frac{C^{\prime}}{8}\right)$ where $C^{\prime}:=\frac{1}{\lambda_{1}}=\frac{1}{648}$, SGD-SLS with $\eta_{\max}=\frac{1}{\epsilon}$ and $c=\frac{2}{3}$ requires
$T=O\left(\min\left\{n,\frac{1}{\epsilon}\right\}\;\frac{1}{\gamma^{2}}\;\left[\ln\left(\frac{n}{\epsilon^{2}}\right)\right]^{2}\right)$
iterations to ensure that $\mathbb{E}[\inf_{t\in[T]}f(\theta_{t})]\leq\frac{5\epsilon}{2}$.
Proof.
Combining Thm. 4 and Thm. 5 completes the proof. ∎
See 6
Proof.
Note that from Lemma 1, we know that ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{C^{\prime}}{{f_{t}}(\theta_{t})}\right\}$ where $C^{\prime}:=\frac{1}{\lambda_{1}}$. Since $f_{t}(\theta)\geq 0$, combining this with Lemma 5, we get that,
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\min\left\{\eta_{\max}{f_{t}}(\theta_{t}),C^{\prime}\right\}+2\eta_{\max}[{f_{t}}(u)]$
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{\eta_{\max}{f_{t}}(\theta_{t}),C^{\prime}\right\}}{2}+2\eta_{\max}[{f_{t}}(u)]$ (Setting $c=\frac{2}{3}$)
Since the algorithm guarantees that $f_{i}(\theta_{t})\geq\frac{\epsilon}{2}$ for all $t$, ${f_{t}}(\theta_{t})\geq\frac{\epsilon}{2}$. Hence,
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{1}{2}\,\min\left\{\frac{\eta_{\max}\,\epsilon}{2},C^{\prime}\right\}+2\eta_{\max}[{f_{t}}(u)]$
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{1}{2}\,\min\left\{\frac{1}{2},C^{\prime}\right\}+\frac{2{f_{t}}(u)}{\epsilon}$ (Setting $\eta_{\max}=\frac{1}{\epsilon}$)
$\gamma$ $:=\min_{i}\,y_{i}\langle x_{i},u^{*}\rangle$ (32)
For a scalar $\beta>0$,
Consider $u=\beta u^{*}$ where $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon^{2}}\right)$ implies that $f(u)\leq\epsilon^{2}$. Since $\epsilon\leq 1$, $f(u)\leq\epsilon$. Using this relation with the above inequality, we get,
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{1}{2}\,\min\left\{\frac{1}{2},C^{\prime}\right\}+2\epsilon$
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{1}{4}\,\min\left\{\frac{1}{2},C^{\prime}\right\}$ (Since $\epsilon\leq\frac{1}{8}\,\min\left\{\frac{1}{2},C^{\prime}\right\}$)
Taking expectation w.r.t. the randomness from iterations $t=0$ to T1T-1,
$\mathbb{E}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\mathbb{E}[\left\|\theta_{t}-u\right\|_{2}^{2}]-\frac{\min\left\{\frac{1}{2},C^{\prime}\right\}}{4}$
$\implies\mathbb{E}[\left\|\theta_{T}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{\min\left\{\frac{1}{2},C^{\prime}\right\}}{4}\,T$ (Summing from $t=0$ to T1T-1)
$\leq\beta^{2}-\underbrace{\frac{\min\left\{\frac{1}{2},C^{\prime}\right\}}{4}}_{:=C}\,T$ (Since $\theta_{0}=0$ and $\left\|u\right\|=\beta$)
$\mathbb{E}[f(\theta_{T})-f(u)]$ $\leq\frac{\epsilon}{2}+\underbrace{[\nu^{2}+1]}_{:=L}\,\mathbb{E}\frac{\left\|\theta_{T}-u\right\|_{2}^{2}}{2}\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left[\beta^{2}-C\,T\right]$
$\implies\mathbb{E}[f(\theta_{T})]$ $\leq\frac{3\epsilon}{2}+\frac{L}{2}\,\left[\beta^{2}-C\,T\right]$ (Since $f(u)\leq\epsilon$)
In order to ensure that $\mathbb{E}[f(\theta_{T})]\leq\frac{3\epsilon}{2}$, it is sufficient to set $T$ as:
$T\geq\frac{\beta^{2}}{C}$ $=\frac{4}{\min\{C^{\prime},\frac{1}{2}\}\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}$